OpenStudy (sugarmochi):

Please please help Find the value of m if the line 3x+4y-5=0 is parallel to the line (m-2)x-2y+2=0

zepdrix (zepdrix):

If the lines are parallel then they have the same slope/gradient value. One idea is to put each equation into slope-intercept form, and match up the slopes from there. Do you understand how to put this into slope-intercept form?\[\large\rm 3x+4y-5=0\qquad\qquad\to\qquad\qquad y=?\]

OpenStudy (sugarmochi):

y=-3x+5/4 right..?

zepdrix (zepdrix):

If you mean (-3x+5)/4 then yes. Brackets are important. Both terms are being divided by the 4.

zepdrix (zepdrix):

So we can clearly see that the slope of the first line is -3/4. How bout the other one?

OpenStudy (sugarmochi):

not sure-

zepdrix (zepdrix):

\[\large\rm (m-2)x-2y+2=0\]We'll subtract 2 to the other side,\[\large\rm (m-2)x-2y=-2\]We'll subtract the x stuff to the other side,\[\large\rm -2y=-(m-2)x-2\]Let's multiply by -1 to get rid of all of the negative stuff,\[\large\rm 2y=(m-2)x+2\]And then divide by 2, ya?

zepdrix (zepdrix):

\[\large\rm y=\frac{m-2}{2}x+\frac{2}{2}\]

zepdrix (zepdrix):

Confused by any of those steps? :o What's the slope of this second line?

OpenStudy (sugarmochi):

ahh i get the steps- would the slope of the second line be \[\frac{ m-2 }{ 2}\] or \[\frac{ -2 }{ 2 }\]

zepdrix (zepdrix):

It's the stuff connected to the x, so the (m-2)/2. Ok great we've found our slopes! Since these lines are parallel, our slopes are equal,\[\large\rm -\frac34=\frac{m-2}{2}\]From there, solve for m.

OpenStudy (sugarmochi):

m=1/2?

zepdrix (zepdrix):

Yayyyy good job \c:/

OpenStudy (sugarmochi):

asdfghjkl thankyou so much ! Oh- is it alright for you to help me with another question?

zepdrix (zepdrix):

Sure, I prolly have time for another c:

OpenStudy (sugarmochi):

Find the equation of the line through the intersection of x-y+3=0 and 2x-3y+5=0 with gradient -3/4

zepdrix (zepdrix):

x- y+3=0 2x-3y+5=0 So we'd like to know where the lines intersect, or in other words, the point where they have the same (x,y) value. We have a couple of options: We can either solve one of the equations for y and plug it into the other, this is called `substitution`. Or we can add/subtract the equations together, this is called `elimination`. Are you familiar with either method?

OpenStudy (sugarmochi):

substitution-?

zepdrix (zepdrix):

Mmm ok we can try that. So solve the first equation for y. y=?

OpenStudy (sugarmochi):

y=x-3?

zepdrix (zepdrix):

Ok good. So the \(\large\rm \color{orangered}{y}\) in our first equation is \(\large\rm \color{orangered}{(x-3)}\). We want these equations to have the same y value, \[\large\rm 2x-3\color{orangered}{y}+5=0\]So we're going to let this y in the second equation be the same y from the first, ya?\[\large\rm 2x-3\color{orangered}{(x-3)}+5=0\]We can solve for x from this point.

OpenStudy (sugarmochi):

x=-4?

zepdrix (zepdrix):

I'm not sure, I didn't work it out just yet >.< 2x-3x+9+5=0 -x+14=0 Hmm -4 that doesn't look quite right. I think maybe you forgot to distribute the -3 to the -3, ya?

zepdrix (zepdrix):

No actually I think you're right :) I made a boo boo somewhere. Lemme see where...

zepdrix (zepdrix):

Oh oh, ok. The first equation should have given us y=(x+3). That was the issue.

zepdrix (zepdrix):

\[\large\rm 2x-3\color{orangered}{(x+3)}+5=0\]

zepdrix (zepdrix):

Which yes, will lead to x=-4.

zepdrix (zepdrix):

Plug that value into the first equation to find a corresponding y value.

OpenStudy (sugarmochi):

y=-1..?

zepdrix (zepdrix):

Ok great! So we've discovered that these lines intersect at \(\large\rm (x_o,y_o)=(-4,-1)\)

zepdrix (zepdrix):

So our new line will pass through this point, and will have a gradient value of -3/4. Are you familiar with point-slope form of a line? \(\large\rm y-y_o=m(x-x_o)\) Because we have all of the necessary pieces to just plug right in.

OpenStudy (sugarmochi):

it would be y+1=-3/4(x+4) ?

zepdrix (zepdrix):

Yayyy good job \c:/